OpenStudy (anonymous):

1/4x+8=5/8x+2 I've simplified to3/8x +8=2 Now I'm stuck

OpenStudy (jack1):

so is that \(\large \frac 14 x + 8 = \frac 58 x + 2\)?

OpenStudy (campbell_st):

well in simplifying I think you have made a mistake... if you subtract 1/4 x from both sides you get \[\frac{3}{8} x + 2 = 8\] next subtract 2 from both sides then multiply by 8 and lastly divide by 3